# Unavoidable Canonical Nonlinearity Induced by Gaussian Measures Discretization

### v4

@@ I Introduction
In the present work, we define the discretization-induced geometric contribution for Gausian families as an unavoidable canonical nonlinearity (UCN), and demonstrate that it admits a consistent information-geometric interpretation, thereby addressing this missing piece. To quantify the UCN, a framework is required that is sensitive to probability weights as well as to the geometric rearrangement of probability mass induced by discretization. Motivated by this requirement, we employ optimal transport theory, specifically the 2-Wasserstein distance $W_{2}$. By introducing a quadratic cost function aligned with the Fisher metric of Gaussian families, the Wasserstein distance becomes directly compatible with information-geometric descriptions while remaining sensitive to discretization-induced geometric distortions. We derive an explicit expression for $W_{2}$ between a continuous Gaussian and its discretized counterpart in the limit of vanishing discretization scale $d\to 0$, yielding a universal form depending only on the covariance matrix of the CDOS and $d$. We further show that this limiting Wasserstein distance admits a clear geometric interpretation on the statistical manifold, equivalent to a KL divergence associated with the expected parallel translations of continuous Gaussian. The present study thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems, which has remained inaccessible to information-geometric approaches alone.
In the final part of this paper, we show that the derived $W_{2}$-KL equivalence can be extended to a broader class of statistical submanifolds beyond Gaussian families, under the condition that the transport cost is aligned with the Fisher metric associated with the relevant degrees of freedom in system of interest. The details are shown below.
II Concept and Derivation
Unavoidable Canonical Nonliearity
@@ II Concept and Derivation
Wesserstein Distance Aligned with Fisher Metric
@@
Under these preparations, we should quantify $W_{2}\left(P_{c},P_{d}\right)$, i.e., 2-Wesserstein distance between a continuous multivariate Gaussian distribution $P_{c}$ with mean vector $\mu$ and covariance matrix $\Gamma$, and its discretized counterpart $P_{d}$. Here, we consider discretization process as partitioning the configuration space ($\simeq\mathbb{R}^{f}$) into hypercubic cells $V_{k}$ of side length $d$, then taking the vanishing discretization limit of $d\to 0$. Explicitly, this process is expressed as
$^{\forall}k,\ P_{d}\left(q_{k}\right)=\int_{V_{k}}P_{c}\left(q\right)dq.$ (7)
@@
with representative hypercubic as $V_{0}$, and taking Tailor series expantion of $P_{c}\left(q^{\prime}_{k}+u\right)$ around $q^{\prime}_{k}$, we can reasonably retain the leading order in $d$, thereby
@@
$\delta\xi=O\left(h\left(d\right)\right),\ h\left(d\right)=a\cdot d,$ (20)
@@
We also comment on the role of parametrization and invariance in the present Wasserstein-KL correspondence. While the KL divergence itself is invariant under reparametrizations on the statistical manifold, Eqs. (24)-(25) suggest that its expectation under $\rho$ depends on the trace of Fisher metric $\Omega$. This does not represent a contradiction, but can be reasonably understood from the following two aspects. (i) Taking the expectation $\mathbb{E}_{\rho}$ in Eq. (25) is an extrinsic operation living outside the statistical manifold, where the infinitesimal change in the selected parameter, $\delta\xi$, is linked to the discretization scale $d$ and subsequently extended to a random vector (the associated probability density $\rho$ typically reflects a modeling choice related to discretization or coarse-graining, depending on the problem under consideration). (ii) In general, when higher-order terms are present, the KL divergence expanded up to second order in $\delta\xi$, yielding a quadratic form, is explicitly governed by the Fisher metric $\Omega$ associated with the chosen parameter $\xi$. In such cases, once combined with the discretization scale and the extrinsic expectation over parameter variations, the resulting numerical value depends on the chosen parameter and Fisher metric. From these viewpoints, the explicit expression $d^{2}\mathrm{Tr}(\Omega)/12$ generally depends on the chosen parameter, yet the correspondence between the 2-Wasserstein transport cost and the expected KL divergence should be understood as an invariant structural relation. The exception is the Gaussian families considered in the UCN analysis, where (i) the Fisher metric is naturally identified through the isomorphism between the parameter $\mu$ on the statistical manifold and the transport coordinate, and (ii) the natural parameterization and the standard quadratic transport cost lead to $a=1$. While this value of $a$ is not mathematically unique, it corresponds to a canonical choice in which KL divergence on the statistical manifold and the transport cost induce the same local quadratic form, so that no additional rescaling is required. In more general settings, different choices for parameters typically lead to different numerical prefactors. Nevertheless, the universal correspondence $W_{2}^{2}\propto 2\cdot\mathbb{E}_{\rho}[D]$ is always preserved, expressing the robust emergence of second-order information-geometric structure from discretization-induced transport costs. The explicit prefactor depends only on the extrinsic identification between the discretization scale and infinitesimal parameter variations.
@@
Under discretization, the lattice scale $d$ and infinitesimal parameter variations $\delta\xi$ capture a common notion of scale dependence: On the statistical manifold, $\delta\xi$ controls the distinguishability of nearby distributions, while on the level of the underlying sample space, $d$ governs the transport cost that measures geometric distortions between their supports. To consistently quantify this scale-dependent correspondence on the transport side, its cost function should encode the geometric information controlling the infinitesimal distinguishability on the statistical manifold, i.e., employing a quadratic cost canonically consistent with the local Fisher metric. From this correspondence, $d$ is not merely a numerical resolution parameter, but is explicitly linked to $\delta\xi$.
@@
In this sense, the present framework offers a natural and minimal route to interpret discretization-induced nonlinearity in terms of information geometry. Possible alternative metrics or higher-order corrections may encode additional geometric structures; nevertheless, under the requirements of differentiability, quadratic approximability, and a controlled discretization limit, the present transport-information-geometric construction emerges as one of the most natural choices.

### v5

@@ I Introduction
In the present work, we define the discretization-induced geometric contribution for Gaussian families as an unavoidable canonical nonlinearity (UCN), and demonstrate that it admits a consistent information-geometric interpretation, thereby addressing this missing piece. To quantify the UCN, a framework is required that is sensitive to probability weights as well as to the geometric rearrangement of probability mass induced by discretization. Motivated by this requirement, we employ optimal transport theory, specifically the 2-Wasserstein distance $W_{2}$. By introducing a quadratic cost function aligned with the Fisher metric of Gaussian families, the Wasserstein distance becomes directly compatible with information-geometric descriptions while remaining sensitive to discretization-induced geometric distortions. We derive an explicit expression for $W_{2}$ between a continuous Gaussian and its discretized counterpart in the limit of vanishing discretization scale $d\to 0$, yielding a universal form depending only on the covariance matrix of the CDOS and $d$. We further show that this limiting Wasserstein distance admits a clear geometric interpretation on the statistical manifold, equivalent to a KL divergence associated with the expected parallel translations of continuous Gaussian. The present study thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems, which has remained inaccessible to information-geometric approaches alone.
In the following part of this paper, we show that the derived $W_{2}$-KL equivalence can be extended to a broader class of statistical submanifolds beyond Gaussian families, under the condition that the transport cost is aligned with the Fisher metric associated with the relevant degrees of freedom in system of interest.
We will see that while the present mathematical derivations are elementary, the underlying conceptual structure relies on nontrivial identifications; readers primarily interested in the interpretation may wish to consult the Conceptual Position in the final part.
II Derivation and Concept
Unavoidable Canonical Nonlinearity
@@ II Derivation and Concept
Wasserstein Distance Aligned with Fisher Metric
@@
Under these preparations, we should quantify $W_{2}\left(P_{c},P_{d}\right)$, i.e., 2-Wasserstein distance between a continuous multivariate Gaussian distribution $P_{c}$ with mean vector $\mu$ and covariance matrix $\Gamma$, and its discretized counterpart $P_{d}$. Here, we consider discretization process as partitioning the configuration space ($\simeq\mathbb{R}^{f}$) into hypercubic cells $V_{k}$ of side length $d$, then taking the vanishing discretization limit of $d\to 0$. Explicitly, this process is expressed as
$^{\forall}k,\ P_{d}\left(q^{\prime}_{k}\right)=\int_{V_{k}}P_{c}\left(q\right)dq.$ (7)
@@
with representative hypercubic as $V_{0}$, and taking Taylor series expantion of $P_{c}\left(q^{\prime}_{k}+u\right)$ around $q^{\prime}_{k}$, we can reasonably retain the leading order in $d$, thereby
@@
A brief numerical check of the derived expression of Eq. (10) against the original integral form of Eq. (8) is provided in Appendix.
@@
$^{\forall}k:\ \delta\xi_{k}=O\left(h\left(d\right)\right),\ h\left(d\right)=a\cdot d,$ (20)
@@
We also comment on the role of parametrization and invariance in the present Wasserstein-KL correspondence. While the KL divergence itself is invariant under reparametrizations on the statistical manifold, Eqs. (24)-(25) suggest that its expectation under $\rho$ depends on the trace of Fisher metric $\Omega$. This does not represent a contradiction, but can be reasonably understood from the following two aspects. (i) Taking the expectation $\mathbb{E}_{\rho}$ in Eq. (25) is an extrinsic operation living outside the statistical manifold, where the infinitesimal change in the selected parameter, $\delta\xi$, is linked to the discretization scale $d$ and subsequently extended to a random vector. (ii) In general, when higher-order terms are present, the KL divergence expanded up to second order in $\delta\xi$, yielding a quadratic form, is explicitly governed by the Fisher metric $\Omega$ associated with the chosen parameter $\xi$. In such cases, once combined with the discretization scale and the extrinsic expectation over parameter variations, the resulting numerical value depends on the chosen parameter and Fisher metric. From these viewpoints, the explicit expression $d^{2}\mathrm{Tr}(\Omega)/12$ generally depends on the chosen parameter, yet the correspondence between the 2-Wasserstein transport cost and the expected KL divergence should be understood as an invariant structural relation. The exception is the Gaussian families considered in the UCN analysis, where (i) the Fisher metric is naturally identified through the isomorphism between the parameter $\mu$ on the statistical manifold and the transport coordinate, and (ii) the natural parameterization and the standard quadratic transport cost lead to $a=1$. While this value of $a$ is not mathematically unique, it corresponds to a canonical choice in which KL divergence on the statistical manifold and the transport cost induce the same local quadratic form, so that no additional rescaling is required. In more general settings, different choices for parameters typically lead to different numerical prefactors. Nevertheless, the universal correspondence $W_{2}^{2}\propto 2\cdot\mathbb{E}_{\rho}[D]$ is always preserved, expressing the robust emergence of second-order information-geometric structure from discretization-induced transport costs. The explicit prefactor depends only on the extrinsic identification between the discretization scale and infinitesimal parameter variations.
@@
Under discretization, the lattice scale $d$ and infinitesimal parameter variations $\delta\xi$ capture a common notion of scale dependence. On the statistical manifold, $\delta\xi$ controls the distinguishability of nearby distributions, while on the level of the underlying sample space, $d$ governs the transport cost measuring geometric distortions between their supports. To consistently quantify this scale-dependent correspondence on the transport side, the cost function should encode the geometric information that controls infinitesimal distinguishability on the statistical manifold, namely by employing a quadratic cost canonically consistent with the local Fisher metric. From this correspondence, $d$ is not merely a discretization parameter, but is explicitly linked to $\delta\xi$. Although the link between $d$ and $\delta\xi$, and the subsequent extension of $\delta\xi$ to a random vector can be merged into a single operation, we deliberately separate them to clarify their conceptual roles.
Within the present transport-information-geometric framework, the probability density $\rho\left(\delta\xi\right)$ is intrinsically constrained only through its second moment, whose scaling $O\left(d^{2}\right)$ encodes the common discretization scale shared by statistical distinguishability and transport cost. Under this constraint, the detailed functional form of $\rho\left(\delta\xi\right)$ remains underdetermined. This residual freedom reflects the fact that higher-order structures are not fixed by the intrinsic geometry alone. Instead, they may be selected once additional, extrinsic physical requirements—such as the microscopic origin of the underlying continuous distributions—are specified.
@@
In these senses, the present framework offers a natural and minimal route to interpret discretization-induced nonlinearity in terms of information geometry. Possible alternative metrics or higher-order corrections may encode additional geometric structures; nevertheless, under the requirements of differentiability, quadratic approximability, and a controlled discretization limit, the present transport-information-geometric construction emerges as one of the most natural choices.
@@ IV Acknowledgement
Appendix: Numerical verification of the analytic expression for $W_{2}$ at finite $d$
Figure 1: Left: $d$-dependence of the squared 2-Wasserstein distance $W_{2}^{2}$ obtained from the analytic expression Eq. (10) (“Analytic”) and from the numerical evaluation of Eq. (8) (“Numerical”). Right: $d$-dependence of the difference between the numerical and analytic values of $W_{2}^{2}$, defined as the numerical value minus the analytic one..
In order to briefly confirm the accuracy of the derived 2-Wasserstein distance, we compare the analytical expression given in Eq. (10) with the original integral expression in Eq. (8) for a simple one-dimensional Gaussian case. We set the standard deviation to $\sigma=1$. The numerical evaluation of Eq. (8) is performed using the trapezoidal rule with uniform discretization. The integration range is restricted to $q\in[-5\sigma,5\sigma]$, and the representative point $q^{\prime}_{k}$ of each discretization cell is chosen as its midpoint. The numerical results shown in Fig. 1 indicate that the derived analytic expression for $W_{2}$ accurately reproduces the original integral form, particularly in the small-$d$ regime.